A spherical lump of lead has a mass of 24.0 kg what is the radius Of this sphere?
1
Expert's answer
2020-01-21T03:32:49-0500
"m=\frac{4}{3}\pi \rho r^3"

"24=\frac{4}{3}\pi (7870)r^3"

"r=0.09\ m"